OpenStudy (anonymous):

Choose the value for x that makes the open sentence true. 7 + 6 • x = 4 to the second power + x to the second power 5 4 3 2

OpenStudy (anonymous):

i did work for them all and got answers but none of them math im stuck :(

OpenStudy (anonymous):

match*

OpenStudy (ivettef365):

are you sure the problem is written correctly ?

OpenStudy (anonymous):

oops hold

OpenStudy (anonymous):

its 4 to the second power and x to the second power

OpenStudy (anonymous):

Thanks!

OpenStudy (austinl):

\(\large{7+6x=4^2+x^2}\)

OpenStudy (austinl):

And that is incorrect.

OpenStudy (ivettef365):

ok try to substitute the x for each number to see which one is true

HanAkoSolo (jamierox4ev3r):

@zkrup please do not provide answers, it is against protocol !! read this: http://openstudy.com/code-of-conduct

OpenStudy (austinl):

Get all of the values onto one side of the equality. \(\large{x^2-6x+9=0}\)

OpenStudy (austinl):

And then you need to factor.

OpenStudy (ivettef365):

first try 5 and you have 7 + 6*5 = 4^2 + 5^2 7 + 30 = 16 + 25 37 = 41 not true try the other numbers

OpenStudy (austinl):

@ivettef365 A better way than guess and check is a mathematical method.

OpenStudy (anonymous):

thanks to everyone